If the width of a certain rectangle is increased by 3 and the length decreased by 3, the area is seen to increase by 6. But if the width is reduced by 5 and the length increased by 3, the area decreases by 90. Find the original dimensions.
Show all the steps.

Answer by richwmiller(17219)

l*w=a
(w+3)*(l-3)=a+3
(w-5)*(l+3)=a-90